OpenStudy (anonymous):

An elevator is moving upward at a uniform speed v0 = 0.40m/s. Suddenly a bolt detaches from the ceiling and starts to fall. How long will it take the bolt fall on the floor of elevator if the height of cabin is h = 2m?

OpenStudy (anonymous):

How did u manage to get it

OpenStudy (anonymous):

\[y=ut+0.5g(t ^{2})\] Assume the bolt to be dropping along the Y-axis,with initial velocity 0.4m/s assume the point of dropping of the bolt to be the origin of the co-ordinate axis,hence y=-2.u=0.4m/s,plug the values into the equation.g=-10m/s2. though i have a bit of doubt with the velocity,and wat role does it play in the sum!!!

OpenStudy (mani_jha):

good work dude, welcome to openstudy!

OpenStudy (anonymous):

thanks bro, check the question and explain the last line of my comment

OpenStudy (anonymous):

will be nice to know. let me try it ur way

OpenStudy (anonymous):

how come i get 0.44s

OpenStudy (anonymous):

did u take "g" as 10m/s2 or 9.8m/s2

OpenStudy (anonymous):

9,8 and i think i may have messe it al up when i got to -5=0,4m/s*t +4,9m/s/s*t^2 how would you continue from there.

OpenStudy (anonymous):

i then devided the whole eq by 0.4

OpenStudy (anonymous):

sorry the first value is -2 above and not -5

OpenStudy (anonymous):

solve the quadratic equation to get the value of t.First multiply the equation by 10(the first equation) -2=0.4t-5t^2 therefore 50t^2-4t-20=0. 25t^2-2t-10=0 now solve it p.s. take the value of g as 10m/s2 for your own convenience. Sorry the value should be 1.43seconds and not 0.6 seconds

OpenStudy (anonymous):

clear your mind and try to visualize the question as it is appearing in front of you.

OpenStudy (anonymous):

Thanks i can apply quadratic formula for that, but doesnt 1.43s seem a long time esp that the lift is also moving up.

OpenStudy (anonymous):

i think its okay as the velocity of the lift is not that high,come on 0.4m/s is a small value..